OpenStudy (anonymous):

0 = 2(2x – 1) (x + 7)

OpenStudy (anonymous):

2x-1=0 or x+7=0 2x=1 or x=-7 x=1/2 or -7

OpenStudy (anonymous):

@Lang

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

ur welcome